you are selling photos of landscapes you've taken. You sell them framed for $15 and unframed for $7 each. If you sold 52 photos for a total of $484, how many framed and unframed did you sell
Answer by JulietG(1812)

If all 52 photos were framed, your total would be $780
Your actual total is $484. 780-484 = 296
The difference between the framed and unframed photos is $8 (15-7).
296 / 8 = 37
37 of your photos sold for $7 for a total of $259
52-37, or 15, or your photos sold for $15 for a total of $225
Total amount sold = 259+225 = $484
.
37 unframed
15 framed